Requirement 2. Suppose plant capacity were only 11,000 windows instead of 12,000 windows each month. The special order must either be taken in full or be rejected completely. Should Award Plus accept the special​ order? Show your calculations. Complete the analysis below to determine if Award Plus should accept the special order under this scenario.

Variable costs that vary with number of units produced

Direct materials

$600,000

Direct manufacturing labour

450,000

Variable costs (for setups, materials handling, quality control, and so on)

that vary with number of batches, 200 batches x $1,500 per batch

300,000

Fixed manufacturing costs

200,000

Fixed marketing costs

50,000

Total costs

$1,600,000

Award plus has just received a special​ one-time-only order for 2,000 windows at $250 per window. Accepting the special order would not affect the​ company's regular business or its fixed costs. Award Plus makes windows for its existing customers in batch sizes of 50 windows ​(200 batches x 50 windows per batc 10,000​windows). The special order requiresAward Plus to make the windows in 40 batches of 50 windows

With One-Time

Only Special Order

Under Reduced

Plant Capacity

11,000 Windows

Revenues

Variable costs:

Direct materials

Direct manufacturing labour

Batch manufacturing costs

Fixed costs:

Fixed manufacturing costs

Fixed marketing costs

Total costs

Operating income
Answer #1
Particulars Without Accepting Special Order Accepting Special Order Increase / (Decrease)
Sales Revenue 30,00,000 32,00,000 2,00,000
Variable Costs
Direct Materials 6,00,000 6,60,000 60,000
Direct Labour 4,50,000 4,95,000 45,000
Variable Costs 300,000 330,000 30,000
Total Variable Costs 13,50,000 14,85,000 1,35,000
Contribution Margin 16,50,000 17,15,000 65,000
Fixed Costs
Fixed Manufacturing Costs 2,00,000 2,00,000 0
Fixed Marketing Costs 50,000 50,000 0
Total Fixed Cost 2,50,000 250,000 0
Operating Income 14,00,000 14,65,000 65,000

The Offer should be accepted as it increases the Operating Income by $ 65,000

Notes and Supporting Work

Contribution Margin = Sales - Total Variable Costs

Operating Income = Contribution Margin - Total Fixed Costs

Calculation for Accepting the Sales Order

On Accepting Special Order Sales to Outside will decrease by 1,000 Units

Sales Revenue = 9,000 Units * $ 300 = $ 27,00,000

Sales Revenue for additional sales of 2,000 Units = 2,000 * $ 250 per Unit = $ 500,000

Total Sales Revenue = 27,00,000 + 5,00,000 = $ 32,00,000

Total Variable Costs

Variable Manufacturing Overhead = 220 Batches * $ 1,500 per Batch = $ 330,000

Direct Labour = $ 45 per Unit * 11,000 Units = $ 495,000

Direct Materials = $ 60 per Unit * 11,000 Units = $ 660,000

Additional Information

Direct Materials Cost per Unit = 600,000 / 10,000 = $ 60 per Unit

Direct Labour Cost per Unit = 450,000 / 10,000 = $ 45 per Unit

Calculation for 10,000 Units without special order has already been done in the question.

Increase / (Decrease) = Values after Acceptance of Special Order - Values without Accepting Special Order

Batches for 11,000 Units = 11,000 / 50 = 220 Batches